# An evaluation of LLMs and Google Translate for translation of selected Indian languages via sentiment and semantic analyses

### v2

@@ 6 Conclusion
We presented a framework for the evaluation of the translation of selected Indian languages including Sanskrit (Bhagavad Gita), Telugu (Maha Prasthanam) and Hindi (Tamas) by LLMs (GPT-3.5, GPT 4o, Gemini) and Google Translate. We used a combination of sentiment and semantic analysis to compare the translated text with translations by human experts. Our findings suggest that while LLMs have made significant progress in translation accuracy, challenges remain in preserving sentiment and semantic integrity, especially in figurative and philosophical contexts.
For Sanskrit, GPT-3.5 was the closest to the human expert’s sentiment distribution, followed by GPT-4o, Gemini, and Google Translate. In Telugu, GPT-3.5 again showed the closest alignment with the human expert, with GPT-4o next, followed by Gemini and Google Translate. For Hindi, GPT-3.5 remained the most similar to the human expert, followed closely by GPT-4o, Gemini, and Google Translate. Overall, GPT-3.5 demonstrated the highest consistency with human sentiment analysis across all three languages.
For Sanskrit-English translation, GPT-4o performed best, followed by GPT-3.5, Gemini, and Google Translate (GT) in that order. In Telugu-English translation, Gemini ranked highest, with GPT-4o closely behind, followed by GT and GPT-3.5. For Hindi-English translation, GPT-4o again performed best, followed by Gemini, GPT-3.5, and GT. Overall, GPT-4o showed the strongest performance across most languages, with Gemini excelling in Telugu-English translation.
@@
The semantic analysis scores indicate that the LLMs provide better results in terms of semantic richness and vocabulary given difficult low-resource languages and philosophical texts such as the Bhagavad Gita, which is also difficult to translate by human experts. The variability in the models’ performance in divergent text styles as well therefore gives opportunities to tailor the models to suit the specific requirements of different languages. Furthermore, we found that LLMs furnish better translations even if capturing the context of the text.
@@
The results of this study could help in the development of more accurate and culturally sensitive translation systems for large language models. Prompting the LLMs with some additional context or any background information based on what text has been used for translating would help enhance the quality of translation, can also be done. Future work can be done with the comparison of these models with other texts and languages with more data.

### v3

@@ VI Conclusion
We presented a framework for the evaluation of the translation of selected Indian languages, including Sanskrit (Bhagavad Gita), Telugu (Maha Prasthanam) and Hindi (Tamas) by LLMs (GPT-3.5, GPT-4o, Gemini) and Google Translate. We used a combination of sentiment and semantic analysis to compare the translated text with translations by human experts. Our findings suggest that while LLMs have made significant progress in translation accuracy, challenges remain in preserving sentiment and semantic integrity, especially in figurative and philosophical contexts.
@@
The sentiment analysis revealed that GPT-based models gave the best performance at maintaining the sentiment polarity when compared to the human expert in the translated texts. The semantic analysis scores indicate that the LLMs provide better translation results when compared to Google Translate in terms of semantic richness and vocabulary. We find that LLMs are generally better at translation in terms of sentiment and semantics when compared to Google Translate. The variability in the models’ performance in divergent text styles as well therefore, gives opportunities to tailor the models to suit the specific requirements of different languages. Furthermore, we found that LLMs furnish better translations even if they capture the context of the text.
@@
The results of this study could help in the development of more accurate and culturally sensitive translation systems for large language models. Prompting the LLMs with some additional context or any background information based on what text has been used for translation would help enhance the quality of translation, and can also be done. Future work can be done with the comparison of these models with other texts and languages with more data.